Can someone please help me to understand what to do with these two problems??? I am lost.
1) Give the variance of the following distribution.
X P(X)
0 0.20
1 0.35
2 0.10
3 0.25
4 0.10
A) 1.83
B) 1.31
C) 1.25
D) 1.71

2) What is the standard deviation of the following probability distribution?
X P(X)
0 0.20
2 0.05
4 0.35
6 0.25
8 0.15
A) 2.6
B) 4.7
C) 5.4
D) 3.9

Answer by stanbon(48568)

Can someone please help me to understand what to do with these two problems??? I am lost.
1) Give the variance of the following distribution.
X.. P(X)..X^2.X^2*P(X)
0.. 0.20..00...0.00
1.. 0.35...1...0.35
2.. 0.10...4...0.40
3.. 0.25...9...2.25
4.. 0.10..16...1.60
Add up the X^2*P(X) to get 4.6
Find the mean: Sum[X*P(X)] = 0.35 + 2*0.1 + 3*.25 + 4*0.1 = 1.7
standard deviation = Sqrt[Sum[X^2*P(X)]-mu^2] = sqrt[4.6-1.7^2]
=sqrt[1.71]
Sample Variance = [E(x^2*P(x))]-u^2]]^2 = [sqrt(1.71)]^2 = 1.71
======================================

A) 1.83
B) 1.31
C) 1.25
D) 1.71
2) What is the standard deviation of the following probability distribution?
Follow the same pattern as above:
Answer: 2.6
-------------------------
X P(X)
0 0.20
2 0.05
4 0.35
6 0.25
8 0.15
-----------
A) 2.6
B) 4.7
C) 5.4
D) 3.9
===========
Cheers,
Stan H.